OpenStudy (anonymous):

Hello, anybody help me find an equation of a line that is perpendicular to 3x-y=-4 and passes through the point (-2,4)

OpenStudy (amistre64):

what do we know about properties of perpendicular lines?

OpenStudy (amistre64):

you will also need to be familiar with the point-slope form of a line

OpenStudy (anonymous):

is the answer y=3x+10

OpenStudy (amistre64):

im not sure if your guessing or not .... how would you demonstrate it?

OpenStudy (amistre64):

3x - y = -4 x + 3y = k (-2) + 3(4) = k 10 = k x+3y = 10, so no